OpenStudy (anonymous):

Find the dy/dx of the following function. y= 2cosx/ 1+sinx PLEASE EXPLAIN.

OpenStudy (amistre64):

product rule

OpenStudy (anonymous):

Do you know the rule of division for differentiating.

OpenStudy (amistre64):

forget division; product rule :)

OpenStudy (anonymous):

I think division since its a fraction .

OpenStudy (anonymous):

wouldnt it be the qoutient rule?

OpenStudy (anonymous):

Im pretty sure you use the qoutient rule for solving, but I got lost can someone help?

OpenStudy (amistre64):

y= 2cosx (1+sinx)^-1 y'= 2cosx' (1+sinx)^-1 + 2cosx (1+sinx)'^-1 = -2sinx (1+sinx)^-1 - 2cosx (1+sinx)^-2 (cos(x))

OpenStudy (amistre64):

yougot lost becasue your USING the quotient rule when the product rule is much much simpler to play with

OpenStudy (anonymous):

sorry their divided by eachother not multiplied :/

OpenStudy (amistre64):

actually, they are multiplied to each other ....

OpenStudy (anonymous):

(2cosx)/(1+sinx) can i use the product rule if they are divided by eachother in the original eqaution?

OpenStudy (amistre64):

\[\frac{2cos(x)}{(1+sin(x))}=\frac{2cos(x)}{1}*\frac{1}{(1+sin(x))}=2cos(x)*(1+sin(x))^{-1}\]

OpenStudy (amistre64):

\[\frac{a}{b}=a*b^{-1}\]

OpenStudy (amistre64):

\[y= 2cosx (1+sinx)^{-1}\] \[y'= 2cosx' (1+sinx)^{-1} + 2cosx (1+sinx)'^{-1}\] \[y' = -2sinx (1+sinx)^{-1} - 2cosx (1+sinx)^{-2} (cos(x))\] \[y' = \left(-2sinx (1+sinx) - 2cosx (cos(x))\right) (1+sinx)^{-2}\] \[y' = \frac{-2sinx (1+sinx) - 2cosx (cos(x))}{(1+sinx)^{2}}\]

OpenStudy (anonymous):

thankyou i got to that point, the answer given is -2/1+sinx how do u get to that?

OpenStudy (amistre64):

the answer "given" is a simplified version of this i beleive; and simplified versions pretty much depend on where you want to stop it at. if you plug this into a computer you get the answer back regardless of how simple youve made the final results :)

OpenStudy (amistre64):

\[y' = \frac{-2sinx (1+sinx) - 2cosx (cos(x))}{(1+sinx)^{2}}\] distribute thru and gather up all the likes \[y' = \frac{-2sinx -2sin^2x - 2cos^2x}{(1+sinx)^{2}}\] factor out the -2 and identify the s^2+c^2=1 id \[y' = \frac{-2(sinx +sin^2x +cos^2x}{(1+sinx)^{2}}\] \[y' = \frac{-2(sinx +1)}{(1+sinx)^{2}}\] cancel out the like terms top to bottom \[y' = \frac{-2}{1+sinx}\] and tada!!